# LeapFrog: Accelerating Multiscale Materials Simulations with Machine Learning

### v2

@@ 2.2 Neural Network
The choice of neural network architecture(s) as well as the loss function that it will train on depend greatly on the problem being tackled. In this section, we describe how the multiscale nature of dendritic solidification informed our choices on these two fronts.
2.2.1 Network Architecture
@@
Requiring a machine learning model to predict the interaction of dendrite cells, the emergence of entire side-branches or dendrite extinctions, etc., from input snapshots tens of thousands of time steps before such events even begin — let alone directly predicting final microstructure — poses an enormous challenge. It would require the neural network model to directly predict emergent behaviour in a highly non-linear system; moreover, unlike a second order transition, microstructure arising in a first order transition such as solidification cannot be renormalized under a simple set of scaling laws. A key challenge in developing any neural network predictor of solidification microstructure evolution over some time frame thus becomes how to correlate the outputs of Eqs. (1)-(3) to compound errors made by the Neural Network (NN), such that the output from the NN remains, at any time, within a statistical ensemble of solutions consistent with the predictions of the original phase field equations.
@@
Given the stochastic component of dendritic evolution, we have developed our NN to act as an aid to the PF time-evolution code, a more feasible and reliable objective. This PF aid function consists of predicting the changes in the system’s fields from a final input time $t_{n}=n\cdot dt$ to a projected time $t_{n+m}=(n+m)\cdot dt$, where $dt$ is the physical time interval and $n,m\in\mathbb{Z}^{+}$ integers. A concrete example, the value $m=1000$ was chosen here as over that time span microstructure in the system has time to evolve meaningfully without at the same time spawning entirely new patterns or behaviours.
@@
We employ Long-Short Term Memory (LSTM) networks, a flavor of Recurrent Neural Networks (RNNs) presented by Hochreiter & Schmidhuber 1997[20], which are directly suited to the task of making time-evolution predictions taking into account multiple time-scales. Given a time-series of a system, its Long Term Memory (LTM) and Short Term Memory (STM) pipelines allow it to separately and respectively track consistent trends in the data (e.g., front advancement, coarsening) as well as more immediate ones (e.g., solute diffusion, side-branch growth). This network architecture has been applied in the phase field domain to spinodal decomposition (de Oca Zapiain, Stewart & Dingreville 2021[14], Hu, Martin & Dingreville 2022[15]), epitaxial growth (Qin et al. 2023, in combination with an attention layer[17]), and simple, single-field descriptions of patters emerging in 1st order phase transitions (Peivaste et al. 2022[16]).
@@
LSTMs require additional modification, however, as they were initially designed for the parsing of low-dimensional vectorized data. It is thus necessary to introduce downsampling and upsampling Convolutional Neural Networks (CNNs) (LeCun, 1998[21]) before and after the LSTM (respectively) as well as a convolutional kernel within it. This integration leverages the translationally invariant learning of CNNs as well as their reduced relative memory requirements to allow the network to process the fairly large systems sizes we tackle in phase field modelling. The result is termed a Convolutional Long-Short Term Memory (CLSTM) network.
@@
Finally, the last ML architecture we incorporate is the U-Net architecture (Ronnenberg, Fischer & Brox 2015[22]). CNNs by themselves were not found to be able to keep track of all the length-scales involved in dendritic solidification. U-Nets are particularly well-equipped for this multiscale context. These networks divide the dataflow within the NN into separate downsampling (encoder) and upsampling (decoder) stages that take the input data down to some smallest resolution and back to its input resolution. By itself, this does not differ from the downsampling/upsampling pipeline achievable with CNNs. However, U-Nets have “skip connections” between encoders and decoders that operate at similar resolutions, which allow for the transmission of fine-grain pattern information that would usually get lost during compression to lower resolutions.
@@
Versions of this network structure have been recently used in combination with various “core” architectures by Dingreville et al. 2023 [23] to evolve systems of physical vapor deposition. Notably, however, an LSTM core is not explored in their implementation. This is a significant incorporation that we found is necessary to accurately identify and emulate the multiple time scales of dendritic solidification. The LSTM core conveys additional capabilities to the overall network and resulting simulation algorithm, which will be discussed in Section 2.3.
@@
A basic instantiation of the network that illustrates the general structure can be seen in Fig. 2, which shows the ingestion of one frame of the system’s evolution at $t_{n}$ that gets processed and compressed by the Encoder block, evolved through the CLSTM, and finally decompressed by the Decoder block to output $\Delta_{m}\{\phi,C\}(\mathbf{x},t_{n})$ that will evolve the system fields to $t_{n+m}$. During the evolution by the CLSTM, the Short Term Memory (STM) and Long Term Memory (LTM) states are generated that will inform the evolution of the next snapshot of the system. The above sequence is repeated for every snapshot of the input data.
@@
The network in Fig. 3 undergoes all the same stages as the one described in Fig. 2, but with the addition of one U-net stage and skip connection. The higher skip connection (“Skip Connection 2”) transmits data to the second Decoder before the dimensionality reduction carried out by “Encoder 1”, and thus contains higher resolution system information. The final network structure involves 5 such U-net and skip connection stages. The final number of stages is a result of the lack of improvement in the results when including additional layers, as well as the increased computational cost involved to train the larger resulting network.
Both the number of layers and the number of snapshots included in an input sample were correlated with significant increases in the memory requirements of the weight gradients to be computed during backpropagation. Due to memory constraints imposed by GPU VRAM capacity (NVIDIA A100 40Gb), the MSCLSTM in our application is trained on 5 snapshot inputs, i.e. $\mathbf{I}_{n}\equiv[t_{n-4m},t_{n-3m},...,t_{n}]$. Despite the quite respectable VRAM capacity of the NVIDIA cards, as well as the significant reduction in memory costs allowed by PyTorch’s Adaptive Mixed Precision package [24], managing the space required to store the network, the samples and especially the gradients from backpropagation remains a significant challenge.
Nonetheless, with this iterative process, we can construct a prediction for the system at $t_{n+m}$ which can then also be used as a new input to produce $\Delta_{m}\{\phi,C\}(\mathbf{x},t_{n+m})$, which can itself be used to produce $\{\phi,C\}(\mathbf{x},t_{n+2m})$ , and so on. Again, due to memory constraints, we use a training algorithm that produces the output set $\mathbf{O}_{n}$ necessary to make three $1000dt$ time span jumps (i.e., $\mathbf{O}_{n}\equiv\Delta_{m}\{\phi,C\}[(\mathbf{x},t_{n-4m}),...,(\mathbf{x},t_{n}),(\mathbf{x},t_{n+m}),(\mathbf{x},t_{n+2m})]$). Larger memory resources would allow larger time span predictions of the field configurations.
@@ 2.2.2 Loss Functions
A simple training algorithm that trains the network with the schema described above using the ubiquitous Mean Squared Error (MSE) loss function falls into some “local minima” in the solution space of network configurations. That is to say, the network will train to emulate the “lowest hanging fruit” to quickly minimize its error metric. In this case: the bulk phase. The network predicting $t=121000dt$ in Fig. 1 will quickly learn to paint the top half as entirely solid and the bottom half as entirely liquid, and ignore the comparatively minuscule complex interfacial patterns — no matter their importance in PF.
A correction for this asymmetry in feature focus can be made through a “feature weighted loss function” $\mathcal{L}_{fw}(\mathbf{x},t)$. The main components of this process are to first identify features that are neglected, construct a binary map $\mathbf{M}$ of their presence in the system, and finally to multiply the MSE loss contribution from those pixels by a factor $\alpha^{\prime}$, which yields,
@@
In addition to the mask illustrated in Fig. 4, another feature emphasis on the concentration profile evolution in areas where the target fields have an order parameter value $\phi>\epsilon$ is incorporated into the training. This corrected for an observed compounding error in the predicted concentration profile of the dendrites, while also coincidentally incentivizing mass conservation. Previously, an explicit penalty for not respecting mass conservation across the system was included, but was later removed since it became redundant.
The predictions generated by a network trained on this feature weighted loss function predicts with high-fidelity predictions. Only some small numerical oscillations in the concentration field remain, which can be easily corrected for using post-processing techniques described in Section 3.3.
As a final note, attempts made to inform the network on the dynamics and physics of the system — such as by including a free energy minimization or power spectra comparison term in the loss — proved ineffective and costly. The large number of operations needed to construct these metrics came at a large memory and wall time cost to, effectively, retread the ground covered by PF modelling that we are trying to leap over. Furthermore, these metrics appeared to have a net negative effect on training when compared to identical training runs using only MSE. It is hypothesized that the resulting gradient tree is too convoluted and complex to offer clear weight modification signals to the optimizer through backpropagation.
@@ 2.3 Adaptive LeapFrog Algorithm
The main algorithm developed for this work is designed to couple the MSCLSTM algorithm presented above with with direct simulation of Eqs. (3) to increase the efficiency of simulating microstructure evolution. Specifically, direct PF simulation is used to compute output data from the initial seed up to the early stages of the dendritic regime, after which this output data is input into the trained MSCLSTM. After a set number of time very rapid jumps by the NN amounting to $N_{ML}=n_{ML}\cdot m\cdot dt$ time steps (or, possibly, after a monitored error metric surpasses a threshold of acceptability), the result of the neural network’s time jumps is handed back to the PF code, which can correct any errors accumulated during the MSCLSTM prediction phase, as well as reintroduce explicit thermal fluctuations compared to the $N_{PF}$ time steps, which were evolved without this component of the phase field dynamics. Following this, the NN then takes over again and the above cycle repeats until the solidification front reaches the end of the simulation system (i.e., sample size). The two time-evolution methods each jump in from where the other left off in what we term a “Leapfrog” manner.
It is noted that the implementation of the hand-off between the PF and ML described above is presented in this work as a proof-of-concept, and as such several key features remain rudimentary and left to be improved in future work. Some of these are discussed next.
@@
On the logging front, as signaled in Section 2.1, the direct PF simulator here is being implemented using Adaptive Mesh Refinement (AMR). The network, however, is trained on systems represented on a uniform mesh. As a result, once the PF simulator is done its pre-determined simulation period, it must map its output onto a uniform mesh representation. Conversely, when the ML predictor is done making its predictions, it must hands its output back to the AMR algorithm at the highest level of mesh refinement, after which the AMR algorithm spends the first few time steps coarsening the mesh where appropriate.
A final feature that can be made more efficient is the practice of regular time-stepping. System snapshots are presently generated in numbers whose interval is sufficient to generate an input for the ML algorithm matching the format of the training samples. This is not strictly necessary, but including more snapshots in the input will be likely be of no benefit, as the network has not trained to track trends on those time scales. It would be possible to use an input sample that is shorter than the training samples, however this would run a risk of not providing enough data on the temporal time scale reuired for the network to construct a complete LTM state.
@@
3 Results
This section demonstrates the LeapFrog algorithm described in Section (2) and quantifies its time-acceleration features. We also review the fidelity with which the results of the LeapFrog algorithm match those of direct phase field simulations.
3.1 Acceleration
Top Left: System Time vs. Wall Time comparison between a Phase Field (PF) simulation using Adaptive Mesh Refinement (AMR) (black) and the LeapFrog algorithm (composite) evolution of the same system in the dendritic regime.
Top Right: Breakdown total Wall Time between PF+AMR, an “Idealized” scenario where the PF+AMR data is always organized in memory (minimizing lookup time), and the LeapFrog Algorithm (broken down by process).
Bottom: Comparison of the resulting system field ($\phi$ and $C$) after being evolved with the PF+AMR and the LeapFrog algorithm introduced in this work.
A first example of the time savings of the LeapFrog algorithm are shown in Fig. 5. The data show the results of an application of LeapFrog with $N_{ML}=5\cdot 1000dt$ and $N_{PF}=2\cdot 1000dt$. It can be directly seen in the top left panel that the overall wall time savings afforded by the LeapFrog algorithm increase the longer the system is evolved (in terms of system, i.e., physical, time). Even with the aforementioned overhead of the machine learning pipeline, the total evolution of the system, past some transient time, up to when the dendritic array reaches the end of the simulation domain is achieved using less than half the wall time.
It is noted that the efficiency of the LeapFrog algorithm is greater than an even “idealized” version of the Adaptive Mesh Refinement (AMR) algorithm (blue segments). By “idealized” we mean a version of the AMR that retains its parallelized asynchronous execution of calculations while spatially adjacent cells in the system are always assured to be adjacent, or at least close by, in memory. In practice, results of computations are written to memory out of order in the standard AMR algorithm, which incurs an asymptotic increase in lookup time necessary to find a cell’s neighbours for gradient computation. An optimal time interval could be determined where the overhead needed to reset the locations of cells in memory would result in an overall speedup. However, for our purposes, this is not done here for the standard AMR algorithm. LeapFrog always benefits from a quasi-ideal performance of the AMR mesh because machine learning predictions are always written to AMR input files in order. This is seen in the top left frame of Fig. 5, which shows that the slope of the AMR phase of the LeapFrog algorithm (data in blue) has a lower slope than the direct AMR algorithm, which as described above is not memory-cache optimized by default.
Average Wall Time per $1000dt$ of System Time
LeapFrog
MSCLSTM $2.24$\mathrm{s}$$
MSCLSTM (With Overhead) $6.86$\mathrm{s}$$
AMR
AMR $46.58$\mathrm{s}$$
Ideal AMR $34.02$\mathrm{s}$$
The overall microstructure evolution does present some differences between direct PF+AMR and LeapFrog-enabled PF algorithms. These mainly involve interface fluctuations affected by the presence of thermal noise. An example of this is seen in a dendrite extinction event between the last two dendrites on the right side of the bottom two frames of Fig. 5. The formation of the halted (extinguished) tip appears slightly delayed between the two approaches, and the subsequent side-branch formation of the surviving dendrite arms is somewhat dampened in the LeapFrog algorithm. We expect that the discrepancy of the two approaches can be mitigated by balancing of $N_{ML}$ to $N_{PF}$ due to the absence of noise in the ML predictions. It is currently unclear how to incorporate to implement noise in the NN weights such as to, say, produce thermal fluctuations at interfaces, let alone satisfy the fluctuation-dissipation theorem.
3.2 Adaptive Time Stepping
To explore the adaptability of the LeapFrog algorithm in trading off between reducing simulation time and fidelity of computed output, another run is performed on the same initial seed as the one that produced the system in Fig. 5. This time, however, the algorithm is used with $N_{ML}=5\cdot 000dt=N_{PF}$. The results are shown in Fig. 6.
Top Left: System Time vs. Wall Time comparison between a Phase Field (PF) simulation using Adaptive Mesh Refinement (AMR) (black) and the LeapFrog algorithm (composite) evolution of the same system in the dendritic regime.
Top Right: Breakdown total Wall Time between PF+AMR, an “Idealized” scenario where the PF+AMR data is always organized in memory (minimizing lookup time), and the LeapFrog Algorithm (broken down by process).
Bottom: Comparison of the resulting system field ($\phi$ and $C$) after being evolved with the PF+AMR and the LeapFrog algorithm introduced in this work.
Going similarly through the panels, despite the less exploitative $N_{ML}/N_{PF}$ ratio, we can see that we continue to obtain over a five-fold increase in simulation speed (per $1000dt$) leading to half the wall time needed to arrive at the end of the simulation domain. At the same time, we observe improved fidelity of side-branching and bulk/interface morphology well behind the solidification front, which demonstrates that the LeapFrog algorithm can be adjusted in its stepping ratio to adapt to the level of fidelity as required for the application.
We have found that the LeapFrog-simulated microstructure for the data of Fig. 5 remains true to the direct PF-simulated counterpart up to $t=301000dt$, which illustrates the network adequately performing well past the temporal bounds of its training set ($t=250000dt$). This suggests that the transformations the network has learned to emulate PF-solidification in a generalized manner, not restricted to a location or time within the system’s evolution.
3.3 Prediction Quality of $\phi$ and $C$ Fields
The first numerical characteristics we examined are profiles of the concentration field (segregation). The main interest is verifying the concentration profiles: (i) along the core (i.e., center) of one of the developed dendrites arms, (ii) along isotherms within the liquid, and (iii) along isotherms within the solid. The bulk concentration values within the core of the dendrites were used to verify that the LeapFrog algorithm produces output consistent with known physics such as the Gibbs Thomson and flux boundary conditions at play at a moving solid-liquid interface, which are satisfied to an excellent degree by the present Phase Field (PF) model [2]. The isotherm profiles, on the other hand, allow us to verify that the thermodynamics that underpin phase diagrams buried within the PF formulation is also respected.
Initially, simply taking the outputs of the LeapFrog (LF) algorithm, we obtain the profiles illustrated in Fig. 7 that exhibit the oscillatory behaviour mentioned in Section 2.2.2.
This is corrected by applying a smoothing filter to the area within the dendrites that were produced by the algorithm. This is implemented by producing a mask analogous to those used in feature weighted loss functions, illustrated in Fig. 8, and applying a unitary smoothing operator at the indicated areas.
b) The magnitude of the gradients in said concentration field.
This outputs concentration profiles that almost exactly match the concentration profile predicted by direct PF simulations, as shown in Fig. 9.
Figure 9: Comparison of concentration profiles along a dendrite’s core (centerline) for a simulation evolved with Adaptive Mesh Refinement (AMR) and LeapFrog (LF) at $t^{*}=250000dt$. (red lines in Fig. 7) with smoothing mask illustrated in Fig. 8 applied to the $C$-field output from the LeapFrog algorithm. The core (centerline) concentration profile is shown on the left. Also shown are the concentration profiles along the two isotherms in the liquid (top right) and solid (bottom right) phases. The point $y=0$ corresponds to the top of the centerline in the left sub-frames.
Another key metric of any directional solidification model is the primary arm spacing (PAS). The PAS is related to the dominant peak in the power spectra of the $\phi$ and $C$ fields. We thus examined the PAS by taking the mean of the power spectrum over 10 randomly initialized seeds. First, the initial seeds are evolved until $t=250000dt$ using only PF. Copies of these systems at $t=65000dt$ were then also evolved using a $N_{ML}=5000dt$, $N_{PF}=2000dt$ LeapFrog configuration. At time $t=250000dt$, field power spectra are extracted from each configuration in the ensemble in both cases. Final average power spectra along with a standard deviation for each $k$-mode are shown in Fig. 10. There is strong agreement between the ensemble power spectra profiles of the two approaches, and thus the predicted primary arm spacings $\lambda_{1}$ match as well. Our results show that, despite not reproducing an identical arrangement of microstructure as the PF-evolved system, LF still evolves the system in a manner that explores the same statistical ensemble consistent with initial conditions.
Figure 10: Mean power spectra comparison with uncertainty intervals between direct Phase Field(PF)/AMR simulations (black) and LeapFrog(LF)-enabled PF simulation (red) for both the phase (left) and concentration (right) fields. Red/Black barred areas designate regions of overlap between the PF and LF uncertainty. The power spectra have been constructed from 10 simulations with different randomly initialized seeds, and the $k=0$ mode set to 0.
4 Discussion
We have presented an algorithm that is capable of accelerating multiscale phase field simulations of dendritic solidification of dilute binary alloys using Adaptive Mesh Refinement (AMR). This is achieved by training a neural network that combines Convolutional Neural Networks, Long-Short Term Memory networks and U-Net architectures to tackle the large system sizes containing numerous time-scales and a large range of length-scales inherent to the microstructure.
We have also demonstrated that the training of the neural network can be focused on features of interest by developing simple masks that highlight the locations where said features are present. These masks proportionally amplify their contribution to the final loss metric. This approach is particularly beneficial because it allows for the targeting of very small or very faint patterns in the microstructure evolution, without resorting to direct calculations based on the fields of the model.
We construct an adaptive time stepping algorithm that uses the above neural network (NN) architecture in tandem with a direct phase field (PF) simulator to generate microstructure predictions. These are combined in a “LeapFrog” (LF) fashion, whereby results from the NN are fed into the PF simulator — and vise versa. This LF algorithm is also versatile, since its utilization of ML and PF predictions can be tuned to either favor simulation acceleration or high fidelity with respect to stochastic and fine-grained features, depending on the needs of the application. In its current prototypical form, where no explicit effort was make to reduce the overhead of passing information back and forth from the NN to the PF simulator, the combined simulation platform can provide a speedup rate 5–10 times faster than direct PF simulation. with this speed up being much larger when overhead is factored out.
The speed of the algorithm is also shown to not come at the cost of accuracy, as its predictions — insofar as microstructure statistic are concerned — exactly match those of traditional PF modelling. These include micro-segregation patterns, and primary spacing selection.
Various future directions and extensions of this work are possible. Most directly, the direct integration of the NN into the PF simulator would further increase computational speedup. Furthermore, training the network over a wider range of process conditions $(G,V)$ would allow for predictability over a larger range of phase space. This would significantly reduce the computational cost of using direct PF simulations to explore regions of interest in materials design space. Finally, the network proposed in this work can be modified to act as a local prediction tool in the so-called “mini-mesh” data structure at the heart of the finite difference scheme of the latest Adaptive Mesh Refinement (AMR) platform[6]. This would allow a more direct integration of our NN into AMR, which would further enhance the experimentally relevant system sizes and times scales accessible to phase field modelling.

### v3

@@ 4.2 Neural Network Architecture
The choice of neural network architecture(s) as well as the loss function that they train on depend greatly on the problem tackled. In this section, we describe how the multiscale nature of dendritic solidification informed our choices on the former.
@@
Requiring a machine learning model to predict the interaction of dendrite cells, the emergence of entire side-branches or dendrite extinctions, etc., from input snapshots tens of thousands of time steps before such events even begin — let alone directly predicting final microstructure — poses an enormous challenge. It would require the neural network model to directly predict emergent behaviour in a highly non-linear system; moreover, unlike a second order transition, microstructure arising in a first order transition such as solidification cannot be renormalized under a simple set of scaling laws.
A key challenge in developing any neural network predictor of solidification microstructure evolution over some time frame thus becomes how to correlate the outputs of Eqs. (1)-(3) to compound errors made by the Neural Network (NN), such that the output from the NN remains, at any time, within a statistical ensemble of solutions consistent with the predictions of the original phase field equations.
@@
Given the highly non-linear and stochastic nature of dendritic evolution, we have developed our NN to act as an aid to the PF time-evolution code, a more feasible and reliable objective. This PF aid function consists of predicting the changes in the system’s fields from an input time $t_{n}=n\cdot dt$ to a projected time $t_{n+m}=(n+m)\cdot dt$, where $dt$ is the physical time interval and $n,m\in\mathbb{Z}^{+}$ integers. As a concrete example, the value $m=1000$ was chosen here as over that time span microstructure in the system has time to evolve meaningfully without at the same time spawning entirely new patterns or behaviours.
@@
We employ Long-Short Term Memory (LSTM) networks, a flavor of Recurrent Neural Networks (RNNs) presented by Hochreiter & Schmidhuber 1997[hochreiter_long_1997], which are directly suited to the task of making time-evolution predictions taking into account multiple time-scales. Given a time-series of a system, its Long Term Memory (LTM) and Short Term Memory (STM) pipelines allow it to separately and respectively track consistent trends in the data (e.g., front advancement, coarsening) as well as more immediate ones (e.g., solute diffusion, side-branch growth). This network architecture has been applied in the phase field domain to spinodal decomposition (de Oca Zapiain, Stewart & Dingreville 2021[montes_de_oca_zapiain_accelerating_2021], Hu, Martin & Dingreville 2022[hu_accelerating_2022]), epitaxial growth (Qin et al. 2023, in combination with an attention layer[qin_grainnn_2023]), and simple, single-field descriptions of patters emerging in 1st order phase transitions (Peivaste et al. 2022[peivaste_machine-learning-based_2022]).
@@
LSTMs require additional modification, however, as they were initially designed for the parsing of low-dimensional vectorized data. It is thus necessary to introduce downsampling and upsampling Convolutional Neural Networks (CNNs) (LeCun, 1998[lecun_gradient-based_1998]) before and after the LSTM (respectively) as well as a convolutional kernel within it. This integration leverages the translationally invariant learning of CNNs as well as their reduced relative memory requirements to allow the network to process the fairly large systems sizes we tackle in phase field modelling. The result is termed a Convolutional Long-Short Term Memory (CLSTM) network.
@@
Finally, the last ML architecture we incorporate is the U-Net architecture (Ronnenberg, Fischer & Brox 2015[ronneberger_u-net_2015]). CNNs by themselves were not found to be able to keep track of all the length-scales involved in dendritic solidification. U-Nets are particularly well-equipped for this multiscale context. These networks divide the dataflow within the NN into separate downsampling (encoder) and upsampling (decoder) stages that take the input data down to some smallest resolution and back to its input resolution. By itself, this does not differ from the downsampling/upsampling pipeline achievable with CNNs. However, U-Nets have “skip connections” between encoders and decoders that operate at similar resolutions, which allow for the transmission of fine-grain pattern information that would usually get lost during compression to lower resolutions.
@@
Versions of this network structure have been recently used in combination with various “core” architectures by Dingreville et al. 2023 [oommen_rethinking_2023] to evolve systems of physical vapor deposition. Notably, however, an LSTM core is not explored in their implementation. This is a significant incorporation that we found is necessary to accurately identify and emulate the multiple time scales of dendritic solidification. The LSTM core conveys additional capabilities to the overall network and resulting simulation algorithm, which will be discussed in Section 4.4.
@@
A basic instantiation of the network that illustrates the general structure can be seen in Fig. 10, which shows the ingestion of one frame of the system’s evolution at $t_{n}$ that gets processed and compressed by the Encoder block, evolved through the CLSTM, and finally decompressed by the Decoder block to output $\Delta_{m}\{\phi,C\}(\mathbf{x},t_{n})$ that will evolve the system fields to $t_{n+m}$. During the evolution by the CLSTM, the Short Term Memory (STM) and Long Term Memory (LTM) states are generated that will inform the evolution of the next snapshot of the system. This sequence is repeated for every snapshot of the input data.
@@
The network in Fig. 11 is an expansion of the network in Fig. 10, with the addition of one U-net stage and skip connection, which we use to allow process of larger data inputs. The higher skip connection (“Skip Connection 2”) transmits data to the second Decoder before the dimensionality reduction carried out by “Encoder 1”, and thus contains higher resolution system information.
The final trained version of the network described in this subsection is composed of 5 U-net and skip connection stages. The NN’s final number of U-Net & skip connection stages is a result of the lack of improvement in the results of training sessions including additional layers, as well as the increased computational cost involved to train a larger network.
Said training sessions were performed following a grid search approach for hyperparameter exploration. This method was made feasible given the high number of constraints imposed on the depth of the network by the longest length-scale present in the problem being tackled (See also Ref. [Pinto_MSC]), as well as those imposed by the interplay between data characteristics and hardware limitations.
Both the number of layers and the number of snapshots included in an input sample were correlated with significant increases in the memory requirements of the weight gradients to be computed during backpropagation. Due to memory constraints imposed by GPU capacity (NVIDIA A100 with 40Gb of VRAM), the MSCLSTM in our application is trained on 5 snapshot inputs, i.e., $\mathbf{I}_{n}\equiv[t_{n-4m},t_{n-3m},...,t_{n}]$. Despite the quite respectable VRAM capacity of the NVIDIA cards, as well as the significant reduction in memory costs allowed by PyTorch’s Adaptive Mixed Precision package [micikevicius_mixed_2018], managing the space required to store the network, the samples, and especially the gradients from backpropagation remains a significant challenge.
The training samples $\mathbf{I}_{n}$ are sampled within the time range $t\in[65k,250k]\cdot dt$ described at the end of Section 4.1 from 11 runs simulated through traditional PF+AMR software, with 3 additional runs used for testing and 1 run for validation. In total, the network requires approximately 7 hours to train (although, this training time allows a generous tail to the loss curve that could very feasibly be reduced). Each simulation used here required 3.3 hours to produce, however this time will not be included in the time comparison between the AMR+PF pipeline and the LeapFrog pipeline. This is because the data produced as part of the training data can double up as data used for the final microstructure characterization of a parameter space exploration.
With this iterative process, we can construct a prediction for the system at $t_{n+m}$ which can then also be used as a new input to produce $\Delta_{m}\{\phi,C\}(\mathbf{x},t_{n+m})$, which can itself be used to produce $\{\phi,C\}(\mathbf{x},t_{n+2m})$ , and so on. Again, due to memory constraints, we use a training algorithm that produces the output set $\mathbf{O}_{n}$ necessary to make three $1000dt$ time span jumps (i.e., $\mathbf{O}_{n}\equiv\Delta_{m}\{\phi,C\}[(\mathbf{x},t_{n-4m}),...,(\mathbf{x},t_{n}),(\mathbf{x},t_{n+m}),(\mathbf{x},t_{n+2m}),(\mathbf{x},t_{n+3m})]$). Larger memory resources would allow larger time span predictions of the field configurations.
@@ 4.3 Loss Functions
A simple training algorithm that trains the network with the schema described above using the ubiquitous Mean Squared Error (MSE) loss function falls into some “local minima” in the solution space of network configurations. That is to say, the network will train to emulate the “lowest hanging fruit” to quickly minimize its error metric — in this case: the bulk phases. The network predicting $t=121000dt$ in Fig. 9 will quickly learn to paint the top half as entirely solid and the bottom half as entirely liquid, ignoring the comparatively minuscule complex interfacial patterns — no matter their importance in PF.
Typically, in ML, the number of training samples containing difficult-to-learn features is increased in the training set such that the conceptual importance of said features is adequately represented. In the case of solidification, however, all samples contain a solidification interface that is orders of magnitude smaller than the microstructurally important features whose dynamics it governs through a coupling with the diffusion field, which typically spans longer length scales. Thus, in large systems, the interface comprises an important system volume too small to properly incentivize the network to pay attention to it during training.
To address this, we develop a dynamic correction for the asymmetry in feature focus through a “feature weighted loss function” $\mathcal{L}_{fw}(\mathbf{x},t)$. The main components of this process are to first identify features that are neglected, construct a binary map $\mathbf{M}$ of their presence in the system, and finally to multiply the MSE loss contribution from those pixels by a factor $\alpha^{\prime}$, which yields,
@@
In addition to the mask illustrated in Fig. 12 being applied to the MSE loss, a SoftDice loss term[milletari_v-net_2016] is added as well as another feature emphasis is on the concentration profile evolution in areas where the target fields have an order parameter value $\phi>0.5$ is incorporated into the training (see Equations 4-8 in Supplementary Note 2). The former has been found to facilitate image segmentation in U-Nets used for medical imaging in the presence of significant noise, and the latter corrected for an observed compounding error in the predicted concentration profile of the dendrites, while also coincidentally incentivizing mass conservation. Previously, an explicit penalty for not respecting mass conservation across the system was included, but was later removed since it became redundant.
A network trained on the feature weighted loss function presented above results in a NN capable of generating high-fidelity predictions. Only some small numerical oscillations in the concentration field remain, which can be easily corrected for using post-processing techniques described in Section 2.3.
In closing this section, it is noted that an attempt was made to inform the network on the dynamics and physics of the system — such as including a power spectra comparison term in the loss. The latter was found to improve the length-scale correspondence between the PF+AMR predictions and the LF predictions (see Supplementary Note 2) by a marginal amount, while increasing the GPU memory utilisation by approximately 4Gb.
The memory usage increase is mainly attributed to the large number of operations per pixel that such physical losses can imply, and Supplementay Note 2 illustrates how using a convenient intermediary to the calculation of the full physical metric can provide similar benefits while eschewing the full calculation of said metric (in this case, simply using the magnitude squared of the FFT of the network outputs).
Given the memory constraints mentioned at the end of Section  4.2, the aforementioned MSE, Soft Dice and solid concentration were kept over the power-spectrum-based loss due to their more sizable impact to the overall training and final implemented algorithm results.
@@ 4.4 Adaptive LeapFrog Algorithm
The main algorithm developed for this work is designed to couple the MSCLSTM algorithm presented above with direct simulation of Eqs. (1)-(3) to increase the efficiency of simulating microstructure evolution. Specifically, direct PF simulation is used to compute output data from the initial seed up to the early stages of the dendritic regime, after which this output data is input into the trained MSCLSTM. After a set number of time very rapid jumps by the NN amounting to $N_{ML}=n_{ML}\cdot m\cdot dt$ time steps (or, possibly, after a monitored error metric surpasses a threshold of acceptability), the result of the neural network’s time jumps is handed back to the PF code, which can correct any errors accumulated during the MSCLSTM prediction phase, as well as reintroduce explicit thermal fluctuations compared to the $N_{PF}$ time steps, which were evolved without this component of the phase field dynamics. Following this, the NN then takes over again and the above cycle repeats until the solidification front reaches the end of the simulation system (i.e., sample size). The two time-evolution methods each jump in from where the other left off in what we term a “Leapfrog” manner.
It is noted that the implementation of the hand-off between the PF and ML described above is presented in this work as a proof-of-concept, and as such several key features remain rudimentary and are being improved for future work. Some of these are discussed next.
@@
On the logging front, as signaled in Section 4.1, the direct PF simulator here is being implemented using Adaptive Mesh Refinement (AMR). The network, however, is trained on systems represented on a uniform mesh. As a result, once the PF simulator is done its pre-determined simulation period, it must map its output onto a uniform mesh representation. Conversely, when the ML predictor is done making its predictions, it must hand its output back to the AMR algorithm at the highest level of mesh refinement, after which the AMR algorithm spends the first few time steps coarsening the mesh where appropriate.
A final feature that can be made more efficient is the practice of regular time-stepping. System snapshots are presently generated in numbers whose interval is sufficient to generate an input for the ML algorithm matching the format of the training samples. This is not strictly necessary, but including more snapshots in the input will be likely be of no benefit, as the network has not trained to track trends on those time scales. It would be possible to use an input sample that is shorter than the training samples, however this would run a risk of not providing enough data on the temporal timescale required for the network to construct a complete LTM state.
@@
Data Availability
The datasets used and/or analyzed during the current study available from the corresponding author (damien.pinto@mail.mcgill.ca) upon reasonable request.
Code Availability
The underlying code for this study and training/validation datasets are not publicly available, but may be made available to qualified researchers on reasonable request from the corresponding author (damien.pinto@mail.mcgill.ca).